During a twenty-minute period, a total of 77 vehicles (cars and trucks) passed through a toll booth. The number of cars that came through is five more than twice the number of trucks that came through. What is the difference between the tens digit in the number of cars and the ones digit in the number of trucks? Write a system of equations for this situation and find its solution.
Answer by ankor@dixie-net.com(22740)

During a twenty-minute period, a total of 77 vehicles (cars and trucks) passed through a toll booth.
c + t = 77
The number of cars that came through is five more than twice the number of trucks that came through.
c = 2t + 5
Replace c in the first equation with (2t+5)
(2t+5) + t = 77
2t + t = 77 - 5
3t = 72
t = 24 trucks
find c
c = 2(24) + 5
c = 53 cars
:
What is the difference between the tens digit in the number of cars and the ones digit in the number of trucks?
The 10's of the cars is 5, the ones in the trucks is 4
The difference, obviously, is 1